THE EXTENDED SPECTROSCOPIC DATABASE ON FORMAMIDE: PARENT, ^{13}C AND DEUTERATED SPECIES UP TO 1 THz

A. S. KUTSENKO, Institute of Radio Astronomy of NASU, Chervonopraporna 4, 61002 Kharkov, Ukraine; \underline{R. A. MOTIYENKO}, L. MARGULÈS, Laboratoire PhLAM, CNRS/Université des Sciences et Technologies de Lille 1, Bât. P5, 59655 Villeneuve d’Ascq Cedex, France; J.-C. GUILLEMIN, Sciences Chimiques de Rennes-Ecole Nationale Supérieure de Chimie de Rennes-CNRS, 35700 Rennes, France.

Formamide (NH_{2}CHO) is the simplest interstellar molecule containing a peptide bond that provides polymerization of amino acids. It is also considered as a precursor of acetamide another molecule containing a peptide bond that has been recently discovered in interstellar medium^{a}. While the rotational spectra of the parent istopic species of formamide were extensively studied up to 500 GHz^{b} only few data are available on its deuterated species. We present the new study of the rotational spectra of all singly deuterated isotopologues of formamide as well as new analysis of the rotational spectra of the parent and ^{13}C isotopic species of formamide in the frequency range up to 1 THz. All the measurements have been performed using the Lille spectrometer based on the solid state sources. In total, about 2500 newly measured transition frequencies have been added to existing dataset on the rotational spectra of formamide and its isotopologues. This work is supported by ANR-08-BLAN-0225, the french Programme National de Physique Chimie du Milieu Interstellaire. A.K. would like to acknowledge the support of the Embassy of France in Ukraine.

aHollis, J.M. et al. Astrophys. J. Letters 643 (2006) L25.

bKryvda, A.V. et al. J. Mol. Spec. 254 (2009) 28.